OpenStudy (anonymous):

(9,-6) and (5,6) find the slope of the given pair, if the line is undefined state so

OpenStudy (anonymous):

slope = (y - y')/ (x - x' ) = -6-6/ 9-5 = -12/3 = -4

OpenStudy (anonymous):

so this is undefinded correct

OpenStudy (anonymous):

no no. undefined is when slope is of the form x/0 were x is any real number.

OpenStudy (anonymous):

just 9 -5 = 4 :)